# Analogue of Chern-Simons invariant in non-metricity gravity and axion cosmology

### v1

@@ I Introduction
Axion was originally proposed to solve the strong CP problem [1]. The QCD axion field $\phi$ couples with the instanton density $F\tilde{F}$, which is a pseudo-scalar and therefore the axion field should also be a pseudo-scalar, which has odd parity, that is, under the spatial reflection, the field changes its signature. Such a pseudo-scalar field also appears in the string theories as an imaginary counterpart of the dilaton field [2]. The axion field is a natural candidate for dark matter, and it might be produced by the misalignment of the axion field [3, 4, 5].
@@
For the overall evolution of the FLRW universe expansion, the pseudo-scalar term cannot contribute, but this term changes the dispersion relations of the left- and right-handed modes of the gravitational wave as in the Chern-Simons gravity [16, 17, 18, 19]. . The axion is a pseudo-scalar by definition. In order to say that the field is pseudo-scalar, we need a coupling with a pseudo-scalar quantity such as the Chern-Simons invariant. If there is no such coupling, the field is just a scalar field and cannot be an axion in the original particle physics sense. In non-metricity gravity, we did not know what the pseudo-scalar quantity corresponding to the Chern-Simons invariant is because the non-metricity scalar $Q$ itself is really a scalar quantity and not a pseudo-scalar. Therefore, as a first step, we invent it by using non-metricity tensors.
@@
We choose the potential of the axion field so that the minimum is given by $\phi=0$. When there is a phase transition of the electroweak theory or a phase transition of grand unified theories (GUTs) associated with the spontaneous breakdown of the gauge symmetries, the pseudo-scalar quantity $R_{\mathcal{Q}}$ has a non-trivial value. Then the coupling of the axion field $\phi$ with $R_{\mathcal{Q}}$ generates the misalignment of the axion field, which generates the axion particle production. If the gradient of potential is small enough, the decrease of the dark matter could be slower than $1/a^{3}$, which may solve the $S_{8}$ problem. Furthermore, the GUT phase transition may induce inflation.
@@ II Brief review of non-metricity gravity
In the following, we use $Q_{\sigma\mu\nu}$ to construct the $f(Q)$ gravity.
@@
In addition $Q_{\sigma}$ and $\tilde{Q}_{\sigma}$ are defined by $Q_{\sigma}\equiv Q^{\ \mu}_{\sigma\ \mu}$ and $\tilde{Q}_{\sigma}=Q^{\mu}_{\ \sigma\mu}$. By using the above definitions, we obtain,
@@ III Pseudoscalar quantity
$dx^{\mu}dx^{\nu}dx^{\rho}dx^{\sigma}A_{\mu\nu\rho\sigma}=$ $-\frac{1}{4!}dx^{\mu}dx^{\nu}dx^{\rho}dx^{\sigma}\epsilon_{\mu\nu\rho\sigma}A_{\mu^{\prime}\nu^{\prime}\rho^{\prime}\sigma^{\prime}}\epsilon^{\mu^{\prime}\nu^{\prime}\rho^{\prime}\sigma^{\prime}}$
@@
The pseudo-scalar quantity $R_{\mathcal{Q}}$ is an analogue of the Chern-Simons invariant $R_{\mathrm{CS}}\equiv\tilde{\epsilon}^{\mu\nu\rho\sigma}R^{\tau}_{\ \lambda\mu\nu}R^{\lambda}_{\ \tau\rho\sigma}$ because both of $R_{\mathcal{Q}}$ and $R_{\mathrm{CS}}$ are pseudo-scalar quantities. In some sense, the structure of $R_{\mathcal{Q}}$ is much simpler than that of $R_{\mathrm{CS}}$ because the order of the derivative of $R_{\mathcal{Q}}$ is two but that of $R_{\mathrm{CS}}$ is four. In the case of the instanton density $F\tilde{F}=\epsilon^{\mu\nu\rho\sigma}\mathrm{tr}F_{\mu\nu}F_{\rho\sigma}$, where $F_{\mu\nu}$ is defined by $F_{\mu\nu}=\sum_{a}F_{\mu\nu}^{a}T_{a}=\sum_{a}\left\{\partial_{\mu}A_{\nu}^{a}-\partial_{\nu}A_{\mu}^{q}-f^{a}_{\ bc}\left(A^{b}_{\mu}A_{\nu}-A^{b}_{\nu}A^{c}_{\mu}\right)\right\}T_{a}$. Here $A_{\mu}^{a}$ is the $a$ component of the gauge field and $T^{a}$ is the representation of non-abelian algebra, which is a matrix and $f^{a}_{\ bc}$’s are structure constants of the gauge group. If we compare Eq. (III) with the instanton density, the metric $g_{\mu\nu}$ in (III) might be identified with the gauge field $A^{a}_{\mu}$.
@@ IV A simple model
We may consider the following model,
@@
Then we start with a simple model with axion field $\phi$, whose action is given by,
@@ V FLRW cosmology
In (26) and (27), the redefinition of $\phi$ is absorbed into the redefinition of $\omega\left(\phi^{2}\right)$ as in [26, 27]. Then one may identify $\phi^{2}$ with the cosmological time $t$, $\phi^{2}=t$ and Eqs. (26) and (27) have the following form,
@@ VI Gravitational wave
In order to make the situation clear, we now consider the gravitational wave propagating along the $z$-direction, that is, $h_{ij}\propto\mathrm{e}^{-i\left(\omega t(t)-k(t)z\right)}$, with an angular frequency $\omega$ and a wavenumber $k$. As the spacetime has a translational invariance under the shifts of $x$, $y$, and $z$ directions, one may assume $k$ is a constant. We now consider that the frequency and wavenumber are much larger than the time-evolution of the universe, that is, $\omega^{2},\,k(t)^{2}\gg H^{2},\,\left|\dot{H}\right|$. Under the assumption, we neglect the derivatives of $k(t)$ and the amplitudes with respect to the time $t$.
@@ VII Axion field misalignment and axion particle production
In order that $R_{\mathcal{Q}}$ has a non-trivial value, the universe cannot be the FLRW universe where the spatial part is isotropic and inhomogeneous. When the symmetry breaking occurs, there appear the bubbles of the true vacua and the universe cannot be isotropic or inhomogeneous. Furthermore, the axion production must occur after inflation, so that the density does not become too small due to the rapid expansion of the universe. Therefore, such a period when the universe is neither isotropic nor inhomogeneous could be the period of the electroweak phase transition. As the electroweak mass scale is about 250  $\mathrm{GeV}$, the energy density of the matter should be $\left(250\,\mathrm{GeV}\right)^{4}\sim 4\times 10^{9}\,\mathrm{GeV}^{4}$. Then the nonmetricity scalar $Q$ could be estimated as $Q\sim 10^{9}\,\mathrm{GeV}^{4}/\left(10^{19}\,\mathrm{GeV}\right)^{2}=10^{-29}\,\mathrm{GeV}^{2}$. Because the mass dimension of $R_{\mathcal{Q}}$ is identical with that of $Q$, we may also estimate $R_{\mathcal{Q}}$ as $R_{\mathcal{Q}}\sim 10^{-29}\,\mathrm{GeV}^{2}$.
@@ VIII $S_{8}$ problem
Here $\beta$ and $\epsilon$ are constant, and we assume $\epsilon$ is positive. Furthermore, $\frac{\epsilon}{\beta^{2}}$ plays the role of a small cosmological constant. When $\phi\sim 0$, the potential $V\left(\phi^{2}\right)$ in (49), we find $V\left(\phi^{2}\right)\sim\frac{1}{2}{m_{\phi}}^{2}\phi^{2}+\frac{\epsilon}{\beta^{2}}$, that is, the potential in (44) with the small cosmological constant $\frac{\epsilon}{\beta^{2}}$ is obtained. On the other hand, when $\left|\phi\right|$ is large enough, $\phi^{2}\gg\frac{1}{{m_{\phi}}^{2}\beta^{2}}$, $V\left(\phi^{2}\right)$ goes to a constant, $V\left(\phi^{2}\right)\to\frac{1}{\beta^{2}}\left(1+\epsilon\right)$, and therefore it behaves like a large cosmological constant. This tells that the energy density of the axion field with axion particles could decrease much more slowly than $1/a^{3}$.
@@ IX Beginning of Inflation
In this scenario, during the inflation, the axion field has a non-trivial value and therefore the differences in the dispersion relations of the left-handed mode and the right-handed mode of the gravitational wave. These differences may affect the structure formation. As long as we consider the model where $\alpha\left(\phi^{2}\right)$ is given by (50), $\alpha\left(\phi^{2}\right)$ could be very small due to the Gaussian factor $\mathrm{e}^{-\frac{{m_{\phi}}^{2}\beta^{2}}{2}\phi^{2}}$, and therefore the difference in the polarisation could be suppressed. Of course, this situation depends on the details of the model.

### v2

@@ I Introduction
Axion was originally proposed to solve the strong CP problem [1]. The QCD axion field $\phi$ couples with the instanton density $F\tilde{F}$, which is a pseudo-scalar and therefore the axion field should also be a pseudo-scalar, which has odd parity, that is, under the spatial reflection, the field changes its signature. Such a pseudo-scalar field also appears in the string theories as an imaginary counterpart of the dilaton field [2]. The axion field is a natural candidate for dark matter, and it might be produced by the misalignment of the axion field [3, 4, 5].111 For pioneer works on axion cosmology, see [6, 7, 8, 9, 10, 11], for examples.
@@
For the overall evolution of the FLRW universe expansion, the pseudo-scalar term cannot contribute, but this term changes the dispersion relations of the left- and right-handed modes of the gravitational wave as in the Chern-Simons gravity [22, 23, 24, 25]. The axion is a pseudo-scalar by definition. In order to say that the field is pseudo-scalar, we need a coupling with a pseudo-scalar quantity such as the Chern-Simons invariant. If there is no such coupling, the field is just a scalar field and cannot be an axion in the original particle physics sense. In non-metricity gravity, we did not know what the pseudo-scalar quantity corresponding to the Chern-Simons invariant is because the non-metricity scalar $Q$ itself is really a scalar quantity and not a pseudo-scalar. Therefore, as a first step, we invent it by using non-metricity tensors.
@@
We choose the potential of the axion field $\phi$ so that the minimum is given by $\phi=0$. When there is a phase transition of the electroweak theory or a phase transition of grand unified theories (GUTs) associated with the spontaneous breakdown of the gauge symmetries, the pseudo-scalar quantity $R_{\mathcal{Q}}$ has a non-trivial value. Then the coupling of the axion field $\phi$ with $R_{\mathcal{Q}}$ generates the misalignment of the axion field, which generates the axion particle production. If the gradient of potential is small enough, the decrease of the dark matter could be slower than $1/a^{3}$, which may solve the $S_{8}$ problem. Furthermore, the GUT phase transition may induce inflation.
@@ II Brief review of non-metricity gravity
In the following, we use $Q_{\sigma\mu\nu}$ to construct the non-metricity gravity.
@@
In addition, $Q_{\sigma}$ and $\tilde{Q}_{\sigma}$ are defined by $Q_{\sigma}\equiv Q^{\ \mu}_{\sigma\ \mu}$ and $\tilde{Q}_{\sigma}=Q^{\mu}_{\ \sigma\mu}$. By using the above definitions, we obtain,
@@ III Pseudoscalar quantity
$dx^{\mu}\wedge dx^{\nu}\wedge dx^{\rho}\wedge dx^{\sigma}A_{\mu\nu\rho\sigma}=$ $-\frac{1}{4!}dx^{\mu}\wedge dx^{\nu}\wedge dx^{\rho}\wedge dx^{\sigma}\epsilon_{\mu\nu\rho\sigma}A_{\mu^{\prime}\nu^{\prime}\rho^{\prime}\sigma^{\prime}}\epsilon^{\mu^{\prime}\nu^{\prime}\rho^{\prime}\sigma^{\prime}}$
@@
The pseudo-scalar quantity $R_{\mathcal{Q}}$ is an analogue of the Chern-Simons invariant $R_{\mathrm{CS}}\equiv\tilde{\epsilon}^{\mu\nu\rho\sigma}R^{\tau}_{\ \lambda\mu\nu}R^{\lambda}_{\ \tau\rho\sigma}$ because both of $R_{\mathcal{Q}}$ and $R_{\mathrm{CS}}$ are pseudo-scalar quantities. In some sense, the structure of $R_{\mathcal{Q}}$ is much simpler than that of $R_{\mathrm{CS}}$ because the order of the derivative of $R_{\mathcal{Q}}$ is two but that of $R_{\mathrm{CS}}$ is four. In the case of the instanton density, we have $F\tilde{F}=\epsilon^{\mu\nu\rho\sigma}\mathrm{tr}F_{\mu\nu}F_{\rho\sigma}$, where $F_{\mu\nu}$ is defined by $F_{\mu\nu}=\sum_{a}F_{\mu\nu}^{a}T_{a}=\sum_{a}\left\{\partial_{\mu}A_{\nu}^{a}-\partial_{\nu}A_{\mu}^{q}-f^{a}_{\ bc}\left(A^{b}_{\mu}A_{\nu}-A^{b}_{\nu}A^{c}_{\mu}\right)\right\}T_{a}$. Here $A_{\mu}^{a}$ is the $a$ component of the gauge field and $T^{a}$ is the representation of non-abelian algebra, which is a matrix and $f^{a}_{\ bc}$’s are structure constants of the gauge group. If we compare Eq. (III) with the instanton density, the metric $g_{\mu\nu}$ in (III) might be identified with the gauge field $A^{a}_{\mu}$.
@@ IV A simple model
We may consider the following model with axion field $\phi$,
@@
Then we start with a simple model whose action is given by,
@@ V FLRW cosmology
In (26) and (27), the redefinition of $\phi$ is absorbed into the redefinition of $\omega\left(\phi^{2}\right)$ as in [32, 33]. Then one may identify $\phi^{2}$ with the cosmological time $t$, $\phi^{2}=t$ and Eqs. (26) and (27) have the following forms,
@@ VI Gravitational wave
In order to make the situation clear, we now consider the gravitational wave propagating along the $z$-direction, that is, $h_{ij}\propto\mathrm{e}^{-i\left(\omega t-k(t)z\right)}$, with an angular frequency $\omega$ and a wavenumber $k$. As the spacetime has a translational invariance under the shifts of $x$, $y$, and $z$ directions, one may assume $k$ does not depend on the spatial coordinates $x$, $y$, and $z$. We now consider that the frequency and wavenumber are much larger than the time-evolution of the universe, that is, $\omega^{2},\,k(t)^{2}\gg H^{2},\,\left|\dot{H}\right|$. Under the assumption, we neglect the derivatives of $k(t)$ and the amplitudes with respect to the time $t$.
@@ VII Axion field misalignment and axion particle production
In order that $R_{\mathcal{Q}}$ has a non-trivial value, the universe cannot be the FLRW universe where the spatial part is isotropic and inhomogeneous. When the symmetry breaking occurs, there appear the bubbles of the true vacua and the universe cannot be isotropic or inhomogeneous. Furthermore, the axion production must occur after inflation, so that the density does not become too small due to the rapid expansion of the universe. Therefore, such a period when the universe is neither isotropic nor inhomogeneous could be the period of the electroweak phase transition. The energy scale of the electroweak phase transition is about 160  $\mathrm{GeV}$, the energy density of the matter should be $\left(160\,\mathrm{GeV}\right)^{4}\sim 10^{9}\,\mathrm{GeV}^{4}$. Then the nonmetricity scalar $Q$ could be estimated as $Q\sim 10^{9}\,\mathrm{GeV}^{4}/\left(10^{19}\,\mathrm{GeV}\right)^{2}=10^{-29}\,\mathrm{GeV}^{2}$. Because the mass dimension of $R_{\mathcal{Q}}$ is identical with that of $Q$, we may also estimate $R_{\mathcal{Q}}$ as $R_{\mathcal{Q}}\sim 10^{-29}\,\mathrm{GeV}^{2}$.
@@
Usually, the axion misalignment is believed to have occurred during the inflationary era because the axion decay constant could be estimated to typically ranges from $10^{9}-10^{12}\,\mathrm{GeV}$ if the origin of the axion is QCD. In our model, the axion does not always originate from QCD; therefore, the axion could be more accurately referred to as so-called axion-like particles (ALPs). Therefore, the energy scale of the misalignment in our model is not constrained as in the QCD axion. Furthermore, the misalignment associated with the electroweak phase transition might relax the $S_{8}$ problem as we discuss in the next section.
@@ VIII $S_{8}$ problem
Here $\beta$ and $\epsilon$ are constant, and we assume $\epsilon$ is positive. Furthermore, $\frac{\epsilon}{\beta^{2}}$ plays the role of a small cosmological constant. When $\phi\sim 0$, the potential $V\left(\phi^{2}\right)$ in (49), we find $V\left(\phi^{2}\right)\sim\frac{1}{2}{m_{\phi}}^{2}\phi^{2}+\frac{\epsilon}{\beta^{2}}$, that is, the potential in (44) with the small cosmological constant $\frac{\epsilon}{\beta^{2}}$ is obtained. On the other hand, when $\left|\phi\right|$ is large enough, $\phi^{2}\gg\frac{1}{{m_{\phi}}^{2}\beta^{2}}$, $V\left(\phi^{2}\right)$ goes to a constant, $V\left(\phi^{2}\right)\to\frac{1}{\beta^{2}}\left(1+\epsilon\right)$, and therefore it behaves like a large cosmological constant. This tells that the energy density of the axion field with an particles could decrease much more slowly than $1/a^{3}$.
@@ IX Beginning of Inflation
In this scenario, during the inflation, the axion field has a non-trivial value and therefore there appear the differences in the dispersion relations of the left-handed mode and the right-handed mode of the gravitational wave. These differences may affect the structure formation. As long as we consider the model where $\alpha\left(\phi^{2}\right)$ is given by (50), $\alpha\left(\phi^{2}\right)$ could be very small due to the Gaussian factor $\mathrm{e}^{-\frac{{m_{\phi}}^{2}\beta^{2}}{2}\phi^{2}}$, and therefore the difference in the polarisation could be suppressed. Of course, this situation depends on the details of the model.
@@ X Summary and conclusion
We now summarise the timeline of the scenario in this paper. First, due to the GUT phase transition, the axion misalignment occurs, which may generate inflation. The inflation ends when the axion goes to the bottom of the potential, and due to the oscillation of the axion field, matter particles could be mainly produced. After that, the axion misalignment could occur again due to the electroweak phase transition. Due to this misalignment, the $S_{8}$ problem might be relaxed, and the axion particles could be produced to be dark matter.